# Image-based Multimodal Models as Intruders: Transferable Multimodal Attacks on Video-based MLLMs

### v2

@@ 5 Conclusion
In this paper, we are the first to explore black-box transferable attacks on V-MLLMs. We conduct a thorough investigation of the limitations of existing methods, revealing that they exhibit lower transferability despite their impressive performance in white-box settings. Our findings underscore the need for specially designed transferable attacks tailored to V-MLLMs. We propose the I2V-MLLM attack, a highly transferable cross-modal attack that leverages the intermediate features of IMMs and perturbation propagation to enhance the transferability of attacks targeting V-MLLMs. We hope our work will inspire further research aimed at evaluating and improving the robustness of V-MLLMs.
Supplementary Material
@@ Appendix A Motivation
To improve the transferability of attacks on V-MLLMs, we first conduct a thorough investigation into the shortcomings of existing methods, as showed in Tab.1 of the main content. Based on the experimental results, we summarize the shortcomings of these existing methods as follows: (1) focusing only on sparse key-frames, (2) lacking generalization in perturbing video features, and (3) failing to integrate multimodal information.
@@
Focusing only on sparse key-frames. The FMM attack [21] is the first white-box attack targeting V-MLLMs, using a flow-based temporal mask to perturb selected key frames and disrupt video and LLM features. As shown in Tab.1 of the main content, however, its adversarial transferability is limited due to differences in key-frame selection between the attack and V-MLLMs. Since V-MLLMs may sample unperturbed frames, we replaced the flow-based mask with direct perturbations on key frames sampled by V-MLLMs, calling this the vanilla attack. The comparison between the FMM and vanilla attacks in Tab.1 of the main content shows a marked improvement in white-box performance after the adjustment. However, the transferability remains limited due to diverse frame-sampling strategies in V-MLLMs. Extending key-frame perturbations to the entire video further improves transferability, as shown in rows 1, 2, 4, and 5 of Tab.1 of the main content. These results emphasize the need to perturb all frames sampled by V-MLLMs for optimal attack performance.
@@
Lacking generalization in perturbing video features. The gains from perturbation propagation are limited due to variations in how V-MLLMs extract video features, which often causes the perturbations to overfit to the features of the surrogate model’s video encoder, which reduces the generalizability of perturbations. Enhancing transferability requires targeting common elements across these features. Thus, we focus on lower-level image features. The I2V attack [39], which perturbs each video frame to disrupt image features, demonstrates that using image models as surrogates can effectively generate adversarial samples for video models. Experimental results in row 3, 4 and 5 of Tab.1 of the main content confirm that targeting image features in video frames significantly improves the transferablity of adversarial samples.
@@
Failing to integrate multimodal information. The I2V attack was initially developed for video classification tasks and does not consider the multimodal interactions between video and text, which are crucial for comprehensive video understanding. While I2V attack achieves improved transferability, its effectiveness in video understanding tasks remains limited. Therefore, we propose using an image-based multimodal model as a surrogate, integrating multimodal interaction information into the process of generating adversarial video samples, which leads to a significant improvement in transferability, as demonstrated in the rows 3 and 6 of Tab.1 of the main content.
@@
In summary, we propose using IMMs as surrogates to generate adversarial video samples that incorporate multimodal interactions. In addition, we introduce a perturbation propagation technique to handle different unknown frame sampling strategies. The I2V-MLLM results in Tab.1 of the main content demonstrate the strong transferability of our method across different V-MLLMs.
@@ B.1 Experiment setting
Surrogate models. We perform our proposed approach on three IMMs: BLIP-2 [20], InstructBLIP [9] and MiniGPT-4 [46]. BLIP-2 employs a novel pre-training strategy that integrates frozen pre-trained image encoders and language models, effectively bridging the modality gap with a lightweight Q-Former. InstructBLIP enhances BLIP-2 by introducing vision-language instruction tuning. It uses a diverse range of datasets and an instruction-aware Q-Former to extract relevant features, improving task performance. MiniGPT-4 aligns a frozen visual encoder with the LLMs, achieving a seamless integration that enhances its multimodal capabilities and demonstrates the potential of such alignments in advancing multimodal understanding.
Target models. Our proposed methods are evaluated on four different V-MLLMs: Chat-UniVi [18], LLaVA-Next-Video [44], VideoChat [22] and Video-LLaMA [42], each with a Vicuna-7B [7] as the LLM. Chat-UniVi uses dynamic visual tokens for uniform representation of images and videos, efficiently capturing spatial details and temporal relationships through a multi-scale framework. LLaVA-Next-Video is an open-source chatbot that enhances large language models by fine-tuning them on multimodal instruction-following data. VideoChat combines video foundation models and LLMs with a learnable neural interface, excelling in spatiotemporal reasoning and event localization. Video-LLaMA uses a multi-modal framework that improves LLMs’ understanding of visual and auditory content in videos, employing a Video Q-former to create video encoders and introducing a video-to-text generation task to learn video-language correspondence.
@@
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini [32], which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods [29] to assess Accuracy (Acc.) and GPT-Score. We use GPT-4o-mini [32] as the evaluation model due to its superior performance in text understanding and its cost efficiency.
Fig. 7 illustrates an example of evaluating the VideoQA task with GPT-4o-mini. An adversarial attack is considered successful if the adversarial video sample causes the V-MLLM to change its response to a question from correct (Accuracy: 1) to incorrect (Accuracy: 0).
@@
Figure 7: An example of using GPT-4o-mini to evaluate Accuracy and GPT Score for the VideoQA task, following the methodology in [29].
B.2 Boosting adversarial transferability
To boost adversarial transferability on VideoQA tasks, we draw inspiration from image classification and explore transfer-based attack techniques. Specifically, we consider the Momentum Iterative (MI) method [10], which incorporates momentum to stabilize and diversify gradient updates, and the Nesterov Iterative (NI) method [25], which leverages Nesterov acceleration to refine gradient smoothing and explore a broader adversarial space.
As shown in Tab. 6, the addition of MI and NI does not significantly improve the transferability of FMM and Vanilla methods; instead, it slightly reduces their performance. In particular, NI greatly decreases the attack effectiveness in white-box scenarios, as it smooths the gradients to improve generalization, which inadvertently compromises precision in white-box scenarios.
Although MI and NI do not enhance the attack performance of methods using V-MLLMs as surrogate models, they improve the performance of I2V and I2V-MLLM methods that leverage image models as surrogates. Specifically, NI increases the ASR of I2V against Chat-UniVi from 30.05% to 39.89%, improves the AASR of I2V from 30% to 34.67%, boosts the ASR of I2V-MLLM against Video-LLaMA from 74.91% to 83.08%, and raises the AASR of I2V-MLLM from 55.48% to 60.67%. With the incorporation of MI and NI, the proposed I2V-MLLM method still achieves a significantly higher AASR compared to other methods, further demonstrating its strong transferability and effectiveness in attacking various V-MLLMs.
B.3 Ablation study
The experiments in this section are conducted on the MSVD-QA validation set, using BLIP-2 as the surrogate model. A higher ASR or AASR reflects better adversarial transferability.
Influence of input text. When designing $\mathcal{L}_{P_{v2t}}$ in Eq. (4) of the main content, we consider two types of text inputs: questions and captions generated from the questions and their answers. As illustrated in Fig. 8, we use GPT-4o-mini to generate caption based on the question and answer. The experiment results are shown in Tab. 7, using captions as input yields a slightly higher AASR compared to using questions. This is because captions contain answer-related information, and the perturbations introduced during the iterations disrupt the semantic information within the answers, making it more challenging for V-MLLMs to provide responses aligned with the ground truth.
To disrupt video-level spatial features, I2V-MLLM generates adversarial perturbations by minimizing the cosine similarity between the original and adversarial spatial features:
FMM Chat-UniVi 16.00* 16.34 13.21 21.47 16.76
LLaVA-NeXT-Video 9.22 20.48* 13.49 21.32 16.13
VideoChat 8.12 15.38 14.62* 20.74 14.72
Video-LLaMA 8.70 18.76 13.84 27.93* 17.31
FMM + MI Chat-UniVi 9.63* 17.55 13.12 21.47 15.44
LLaVA-NeXT-Video 8.76 21.72* 13.37 22.22 16.52
VideoChat 7.76 15.86 14.97* 20.74 14.83
Video-LLaMA 8.76 19.75 13.31 28.74* 17.64
FMM + NI Chat-UniVi 13.39* 17.55 13.24 21.61 16.45
LLaVA-NeXT-Video 8.88 20.80* 13.62 23.06 16.59
VideoChat 7.71 15.73 13.37* 19.99 14.2
Video-LLaMA 8.29 19.39 13.43 22.13* 15.81
Vanilla Chat-UniVi 52.48* 19.65 13.71 22.65 27.12
LLaVA-NeXT-Video 9.63 36.85* 13.71 26.51 21.67
VideoChat 8.22 15.86 35.03* 19.81 19.73
Video-LLaMA 11.93 31.82 14.40 63.96* 30.53
Vanilla + MI Chat-UniVi 45.26* 21.05 13.99 22.91 25.05
LLaVA-NeXT-Video 8.96 45.44* 13.71 27.87 23.50
VideoChat 7.55 15.99 40.36* 21.12 21.51
Video-LLaMA 11.78 43.06 13.24 72.22* 35.57
Vanilla + NI Chat-UniVi 14.34* 19.14 13.46 21.23 17.04
LLaVA-NeXT-Video 8.45 23.54* 13.31 24.71 17.75
VideoChat 7.09 15.83 16.88* 19.90 14.94
Video-LLaMA 8.86 26.72 12.87 28.53* 19.50
I2V 30.05 35.62 18.83 30.16 30.17
I2V + MI CLIP-L/14 25.88 34.97 19.63 30.39 27.72
I2V + NI 39.89 39.46 24.48 34.85 34.67
I2V-MLLM 43.39 40.54 63.09 74.91 55.48
I2V-MLLM + MI BLIP-2 48.39 54.71 62.18 73.75 59.76
I2V-MLLM + NI 43.52 49.36 66.72 83.08 60.67
Table 6: ASR(%) of different attacks on the MSVD-QA for VideoQA tasks. AASR represents the average ASR across all target models for each surrogate model. * indicates white-box attacks. A higher ASR or AASR indicates better adversarial transferability. MI [10] refers to the Momentum Iterative method, which enhances gradient diversity by incorporating momentum. NI [25] denotes the Nesterov Iterative method, which improves transferability by smoothing gradients and exploring broader search spaces. The highest attack performance for each target model is shown in bold, and the second-highest in underline.
Target Model Caption Question
Chat-UniVi 0.45 0.44
LLaVA-NeXT-Video 0.36 0.35
VideoChat 0.51 0.50
Video-LLaMA 0.65 0.65
AASR 0.49 0.48
Results on video understanding tasks. Video understanding tasks assess whether V-MLLMs have comprehended the content of a video by posing a range of questions about it. Following Maaz et al. [29], we use a subset of the ActivityNet-200 [15] dataset and employ GPT-4o-mini to evaluate the model’s responses to adversarial examples from five perspectives: Correctness, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency. We compare our proposed I2V-MLLM attack with the Vanilla attack on four V-MLLMs, using clean samples as a reference. Evaluations are performed on white-box Vanilla attack and I2V-MLLM attack (using BLIP-2 as a surrogate model).

### v3

@@ 5 Conclusion
In this paper, we are the first to explore black-box transferable attacks on V-MLLMs. We conduct a thorough investigation of the limitations of existing methods, revealing that they exhibit lower transferability despite their impressive performance in white-box settings. Our findings underscore the need for specially designed transferable attacks tailored to V-MLLMs. We propose the I2V-MLLM attack, a highly transferable cross-modal attack that leverages the intermediate features of I-MLLMs and perturbation propagation to enhance the transferability of attacks targeting V-MLLMs. We hope our work will inspire further research aimed at evaluating and improving the robustness of V-MLLMs.
Acknowledgments
Feng Zheng was supported by the National Natural Science Foundation of China (K23271004). Xi Xiao and Linhao Huang were supported by the Natural Science Foundation of Guangdong Province (2025A1515011946) and the Major Key Project of PCL (PCL2023A06-4). XJ and BH were supported by the NSFC General Program (Project No. 62376235) and the Guangdong Basic and Applied Basic Research Foundation (Project Nos. 2022A1515011652 and 2024A151501239).
Transferability of Adversarial Attacks in Video-based MLLMs:
A Cross-modal Image-to-Video Approach
APPENDIX
@@ Appendix A Motivation
To improve the transferability of attacks on V-MLLMs, we first conduct a thorough investigation into the shortcomings of existing methods, as showed in Tab. 1 in the main content. Based on the experimental results, we summarize the shortcomings of these existing methods as follows: (1) focusing only on sparse key-frames, (2) lacking generalization in perturbing video features, and (3) failing to integrate multimodal information.
Focusing only on sparse key-frames. FMM attack exhibits limited performance when the key-frame ratio is low. This is because FMM selects key-frames based on optical flow and only perturbs these frames, while V-MLLMs typically sample frames sequentially, making it difficult to ensure that all frames extracted by the target model are perturbed.
To validate the impact of the mismatch between FMM’s key-frame selection and V-MLLMs’ sampling strategies on the attack performance, we conduct an ablation study on the key-frame ratio in FMM attack. We perform a white-box FMM attack on Video-LLaMA. As demonstrated in Tab. 6, as the key-frame ratio in FMM increases, the number of adversarial frames sampled by Video-LLaMA also increases, leading to a corresponding improvement in ASR. This demonstrates that the method of selecting key-frames using optical flow in FMM is mismatched with the key-frame selection approach of V-MLLMs, and confirms that injecting perturbations into all frames that V-MLLMs will sample maximizes the attack effectiveness.
Key frame ratio 10 30 50 70 90 100
Sampled frames 1.35 2.67 4.02 5.87 7.36 8.00
ASR 23.12 28.45 34.31 41.02 58.37 64.14
Table 6: The ablation study on the key-frame ratio (%) in the FMM attack is conducted to evaluate its impact on the attack performance. Sampled frames refer to the average number of adversarial frames that Video-LLaMA selects during its forward inference process for each video in the MSVD-QA validation set. (with a maximum of 8 frames).
@@
We replaced the flow-based mask with direct perturbations on all frames sampled by V-MLLMs, calling this the Vanilla attack. The comparison between the FMM and Vanilla attacks in Tab. 1 in the main content shows a marked improvement in white-box performance after the adjustment. However, the transferability remains limited due to diverse frame-sampling strategies in V-MLLMs. Extending key-frame perturbations to the entire video further improves transferability, as shown in rows 1, 2, 4, and 5 of Tab. 1 in the main content. These results emphasize the need to perturb all frames sampled by V-MLLMs for optimal attack performance.
@@
Lacking generalization in perturbing video features. The gains from perturbation propagation are limited due to variations in how V-MLLMs extract video features, which often causes the perturbations to overfit to the features of the surrogate model’s video encoder, which reduces the generalizability of perturbations. Enhancing transferability requires targeting common elements across these features. Thus, we focus on lower-level image features. The I2V attack (wei2021crossmodaltransferableadversarialattacks), which perturbs each video frame to disrupt image features, demonstrates that using image models as surrogates can effectively generate adversarial samples for video models. Experimental results in row 3, 4 and 5 of Tab. 1 in the main content confirm that targeting image features in video frames significantly improves the transferablity of adversarial video samples.
@@
Failing to integrate multimodal information. The I2V attack was initially developed for video classification tasks and does not consider the multimodal interactions between video and text, which are crucial for comprehensive video understanding. While I2V attack achieves improved transferability, its effectiveness in video understanding tasks remains limited. Therefore, we propose using an image-based multimodal model as a surrogate, integrating multimodal interaction information into the process of generating adversarial video samples, which leads to a significant improvement in transferability, as demonstrated in the rows 3 and 6 of Tab. 1 in the main content.
@@
In summary, we propose using I-MLLMs as surrogates to generate adversarial video samples that incorporate multimodal interactions. In addition, we introduce a perturbation propagation technique to handle different unknown frame sampling strategies. The I2V-MLLM results in Tab. 1 in the main content demonstrate the strong transferability of our method across different V-MLLMs.
@@ B.1 Experiment setting
Models. We perform our proposed approach on five I-MLLMs: BLIP-2 (li2023blip2bootstrappinglanguageimagepretraining), InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels), MiniGPT-4 (zhu2023minigpt4enhancingvisionlanguageunderstanding), LLaVA (liu2023visualinstructiontuning) and mPLUG-owl (ye2023mplug). Our proposed methods are evaluated on four different V-MLLMs: Chat-UniVi (jin2024chatuniviunifiedvisualrepresentation), LLaVA-Next-Video (zhang2024videoinstructiontuningsynthetic), VideoChat (li2024videochatchatcentricvideounderstanding) and Video-LLaMA (zhang2023videollamainstructiontunedaudiovisuallanguage), each with a Vicuna-7B (vicuna2023) as the LLM.
Propagation methods. Direct Propagation (DP), Optical Flow-based Propagation (dosovitskiy2015flownet) (OFP), and Bidirectional Linear Interpolation Propagation (dai2017deformable) (BP) are three different perturbation propagation methods. In DP, the adversarial perturbation is directly applied to each corresponding frame segment without any adjustments, simply propagating the perturbation along the frame sequence segments. OFP calculates the optical flow changes between adjacent frames and dynamically adjusts and propagates the adversarial perturbation based on the motion information between frames, facilitating a perturbation transfer that better aligns with frame-to-frame motion patterns. BP employs bidirectional linear interpolation to propagate perturbations from adjacent key frames to the intermediate frames, thereby smoothly transmitting adversarial perturbations across the entire frame sequence.
@@
Metrics. We use Attack Success Rate (ASR) to evaluate the effectiveness of adversarial examples on VideoQA tasks. It measures the percentage of successful attacks on questions the model answered correctly for clean videos. Answer correctness is evaluated using GPT-4o-mini (openai2023gpt4omini), which checks whether the model’s prediction semantically aligns with the ground truth. We also provide the average ASR (AASR) across all evaluated V-MLLMs. A higher ASR or AASR indicates better adversarial transferability. To evaluate the model’s overall performance when encountering adversarial videos, we further employ GPT-assisted methods (maaz2024videochatgptdetailedvideounderstanding) to assess Accuracy (Acc.) and GPT-Score. We use GPT-4o-mini (openai2023gpt4omini) as the evaluation model due to its superior performance in text understanding and its cost efficiency. Fig. 5 illustrates an example of evaluating the Zero-Shot VideoQA task with GPT-4o-mini. An adversarial attack is considered successful if the adversarial video sample causes the V-MLLM to change its response to a question from correct (Accuracy: 1) to incorrect (Accuracy: 0).
@@
Figure 5: An example of using GPT-4o-mini to evaluate Accuracy and GPT Score for the VideoQA task, following the methodology in (maaz2024videochatgptdetailedvideounderstanding).
Since traditional video captioning metrics (Wang_2018_CVPR) such as BLEU, CIDEr, SPICE are not well-suited for evaluating the detailed video captions generated by V-MLLM, we employ GPT-4o-mini as the evaluation model and use the Captioning Score to assess V-MLLM’s performance on the video captioning task. The Captioning Score measures the quality of the model’s predictions by assigning a relative score on a scale from 0 to 100 (see Fig. 6). A higher Captioning Score indicates that the generated captions are closer to the reference captions.
Figure 6: An example of using GPT-4o-mini to evaluate Caption Score for the video captioning task.
B.2 Cross-task transferability
Multimodal interactions and alignments are the core components of multimodal learning, regardless of the specific task. Therefore, we conduct extensive experiments to investigate the cross-task transferability of I2V-MLLM from VideoQA to video captioning tasks. Since traditional video captioning metrics (Wang_2018_CVPR) such as BLEU, CIDEr, SPICE are not well-suited for evaluating the detailed video captions generated by V-MLLMs, we employ GPT-4o-mini as the evaluation model and use the Captioning Score to assess V-MLLM’s performance on the video captioning task. The Captioning Score measures the quality of the model’s predictions by assigning a relative score on a scale from 0 to 100. A higher Captioning Score indicates that the generated captions are closer to the reference captions. We directly transfer the adversarial samples generated from the VideoQA on the MSVD-QA dataset to the video captioning tasks on the MSVD dataset. As shown in Tab. 7, I2V-MLLM causes the largest decrease in the Captioning Scores across all target models, achieving the best attack performance, even surpassing Vanilla and FMM attack methods in the white-box setting. The white-box Vanilla attack, which achieves optimal performance in the VideoQA tasks, performs sub-optimally here because it is specifically designed for the VideoQA task, resulting in slightly weaker cross-task transferability. This highlights that I2V-MLLM not only exhibits cross-model transferability but also demonstrates strong cross-task transferability.
Attack Chat-UniVi LLaVA-NEXT-Video VideoChat Video-LLaMA
Clean 59.64 61.95 54.16 51.60
[…]
Table 7: Captioning Scores on the MSVD dataset for video captioning tasks, ranging from 0 to 100. FMM and Vanilla are white-box attacks (marked with *) on different V-MLLMs. I2V and I2V-MLLM are black-box attacks, with their surrogate models being CLIP-L/14 and BLIP-2, respectively. $\downarrow$ represents the performance drop compared to the clean video samples. The highest attack performance is highlighted in bold.
B.3 Collaboration with textual attacks
In real-world scenarios, attackers often employ a wide range of strategies to attack V-MLLMs, potentially introducing both adversarial video samples and adversarial text into the target models simultaneously. To account for this more practical and comprehensive attack scenario, we conducted additional experiments. Specifically, we applied textual attacks on the MSVD-QA dataset using the Bert-Attack method (li2020bert), modifying only a single character in each question to maximally disrupt the semantic alignment between the video and text.
As illustrated in Tab. 8, Bert-Attack alone achieved an AASR of 37.11%, when combined with I2V-MLLM, the AASR significantly increased to 76%. Notably, Bert-Attack demonstrated stronger performance on Chat-UniVi and LLaVA-NEXT-Video, while our method excelled on VideoChat and Video-LLaMA. The integration of both approaches lead to consistently high ASRs across all target models, highlighting the complementary nature of textual and video-based adversarial perturbations in multimodal attack scenarios.
These findings demonstrate that I2V-MLLM can be seamlessly integrated with textual attack methods to achieve exceptionally robust attack performance, underscoring its substantial disruptive potential in real-world applications.
Attack Bert-Attack (B-A) I2V-MLLM I2V-MLLM + B-A
Chat-UniVi 40.00 48.39 71.13
LLaVA-NEXT-Video 41.62 45.54 66.49
VideoChat 30.89 63.09 79.39
Video-LLaMA 35.92 74.91 87.51
AASR 37.11 57.98 76.31
Table 8: Attack success rates (ASR, %) on the MSVD-QA validation set for Zero-Shot VideoQA tasks. Bert-Attack from (li2020bert).
@@ B.3 Collaboration with textual attacks
B.4 Ablation study
The experiments in following sections are conducted on the MSVD-QA validation set, using BLIP-2 as the surrogate model. A higher ASR or AASR reflects better adversarial transferability.
To disrupt video-level spatial features, I2V-MLLM generates perturbations by minimizing the cosine similarity between the original and adversarial spatial features:
Influence of input text. When designing $\mathcal{L}_{P_{v2t}}$ in Sec. 3.3.2, Eq. (4) of the main content, we consider two types of text inputs: questions and captions generated from the questions and their answers. As illustrated in Fig. 10, we use GPT-4o-mini to generate caption based on the question and answer. The experiment results are shown in Tab. 10, using captions as input yields a slightly higher AASR compared to using questions. This is because captions contain answer-related information, and the perturbations introduced during the iterations disrupt the semantic information within the answers, making it more challenging for V-MLLMs to provide responses aligned with the ground truth.
Target Model Caption Question
Chat-UniVi 45.11 43.68
LLaVA-NeXT-Video 35.67 35.06
VideoChat 51.10 50.10
Video-LLaMA 65.57 64.54
AASR 49.36 48.34
Influence of perturbation bound. To further investigate the impact of perturbation magnitude on attack performance, we tested the ASR under different perturbation bounds ($\epsilon$), including 2, 4, 8, 16, and 32. As shown Tab. 11, the ASR steadily increases with higher perturbation bounds, reaching its peak when the perturbation bound is set to 32. As illustrated in Figs. 9 and 10, with the increase in perturbation bound, the perturbations become more pronounced. To balance attack performance and video vividness, we selected a perturbation bound of 16, which is also a common choice in traditional video attack methods (wei2021crossmodaltransferableadversarialattacks; kim2023breakingtemporalconsistencygenerating; wang2023global).
$\epsilon$ 2 4 8 16 32
Chat-UniVi 12.96 22.76 32.35 48.39 61.32
LLaVA-NeXT-Video 8.14 13.37 24.75 45.54 58.37
VideoChat 25.65 32.14 48.31 63.09 75.23
Video-LLaMA 26.13 34.27 57.24 74.91 84.22
AASR 18.22 25.64 40.66 57.98 69.79
Table 11: Attack success rate (ASR, %) under different perturbation bounds. AASR represents the average ASR across all target models for each surrogate model. A higher AASR indicates better adversarial transferability. The highest attack performance for each target model is shown in bold and the second-highest in underline.
B.5 Results on multi-faceted video understanding tasks.
Multi-faceted video understanding tasks assess whether V-MLLMs have comprehended the content of a video by posing a range of questions about it. Following Maaz et al. (maaz2024videochatgptdetailedvideounderstanding), we use a subset of the ActivityNet-200 (Heilbron_Escorcia_Ghanem_Niebles_2015) dataset and employ GPT-4o-mini to evaluate the model’s responses to adversarial examples from five perspectives: Correctness, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency. We compare our proposed I2V-MLLM attack with the Vanilla attack on four V-MLLMs, using clean samples as a reference. Evaluations are performed on white-box Vanilla attack and I2V-MLLM attack (using BLIP-2 as a surrogate model).